[Ncep.nhc.nco_contacts] TSB Morning Rounds - Thu Nov 12, 2015

Mark DeMaria - NOAA Federal mark.demaria at noaa.gov
Thu Nov 12 13:38:33 UTC 2015

Here is today's update on NHC's computer operations.




The problem Mike reported on the ATCF with pulling data from Opah was only
an issue at one time period. Thus, it may have been due to a temporary
network or other issue on Opah. Although not directly related, Monica is
doing some cleanup and re-organization on Opah to make the data access and
troubleshooting faster in the future.


Jeff will work with Evelyn to provide provide guidance on how to address
the recent text formatter problems.

The AWIPS 15.1.1 upgrade on Tuesday went fairly smoothly. Chris and Matt
are cleaning up a few loose ends.

Two TAFB workstations (Crappie and Cumulus) had nmap/nawips issues over the
past few days. These were fixed by rebooting or logging and logging out.
Matt will run the systems health script to see how long all the operational
NAWIPS systems have been up, and work with TAFB and HSU to reboot those
that are due for reboot.

Other issues:

Two Harris contractors are at NHC today replacing the motors on the new GRB
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20151112/781b6df3/attachment.html 

More information about the Ncep.nhc.nco_contacts mailing list